SOLVENT-MEDIATED ELECTRON LEAPFROGGING: CHARGE TRANSFER IN IBr^{−}(CO_{2}) PHOTODIS-SOCIATION

\underline{SAMANTHA HORVATH}, and ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210; LEONID SHEPS, ELISA M. MILLER, MATTHEW A. THOMPSON, ROBERT PARSON, and W. CARL LINEBERGER, JILA, Department of Chemistry and Biochemistry, University of Colorado at Boulder, Boulder, CO 80309.

In this work, we investigate the time-resolved photoelectron spectra of IBr^{−}(CO_{2}).^{a} In the photodetachment studies performed by Lineberger and co-workers,^{a} IBr^{−}(CO_{2}) is prepared in its electronic ground state (^{2}Σ_{1/2}^{+}) whereupon it is excited to its Ã′ (^{2}Π_{3/2}) excited state, before electron photodetachment/photoionization and dissociation on the C̃ (^{1}Π_{1}) excited state of IBr. Previous experimental work showed that dissociation of bare IBr^{−} yields only I^{−} + Br products.^{b} However in IBr^{−}(CO_{2}), a small fraction (∼ 3%) of the dissociating molecules undergo an electron transfer from I to Br at 350 fs after the initial excitation. Thus a single solvent molecule can initiate a non-adiabatic transition from the Ã′ state to either the lower Ã or X̃ state, thereby producing I + Br^{−} (+ CO_{2}) prior to photoionization. To study the dynamics, we perform high level ab initio calculations (MR-SO-CISD/aug-cc-pVTZ(-PP)) as well as classical molecular dynamics (MD) simulations. The MD simulations capture much of the dynamics of the photodissociation but underestimate the charge-transfer channel. Results of the ab initio calculations show how CO_{2} bend vibrational excitation could increase the percentage of non-adiabatic transitions and how the CO_{2} modifies the charge distribution of IBr^{−} to make the charge transfer accessible. The proposed mechanism and timescales are consistent with the observed Br^{−} products.

aL. Sheps, E. M. Miller, S. Horvath, M. A. Thompson, R. Parson, A. B. McCoy, and W. C. Lineberger, Science, 2010, in press.

bR. Mabbs, K. Pichugin, and A. Sanov, J. Chem. Phys., 2005, 122, 174305.